OpenStudy (anonymous):

Solve 5(1 – 3m) = 30 – 2(4m + 7)

OpenStudy (anonymous):

5-15m=30-8m+7 5-15m=37-8m -5 -5 -15m=32-8m +8m +8m -7m=32 m=-4.5...

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

Hmm, that's not what I got: \[\Large \begin{array}{l} 5(1 - 3m) = 30 - 2(4m + 7)\\ 5 - 15m = 30 - 8m - 14\\ - 15m + 8m = 30 - 14 - 5\\ - 7m = 11\\ m = - \frac{{11}}{7} \end{array}\]